OpenStudy (anonymous):

Please help and show all work, step by step Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

The slope of y = 4x + 7 is 4 or 4/1 The slope of a line perpendicular to y = 4x + 7 must be -1/4 y - (-3) = -1/4(x - 2) y + 3 = -1/4(x) + 1/2 y = -1/4(x) - 5/2 1/4(x) + y = - 5/2 x + 4y = -10